# The role of synthetic data in Multilingual, Multi-cultural AI systems: Lessons from Indic Languages

### v1

@@ 5.2 Data Quality Assessment for Generative Data
Inter-Annotator Agreement: To quantify the reliability of automated evaluation, we computed the percentage agreement between the LLM-judge and the human evaluators. We find significant variance in agreement across metrics (Figure 1). Agreement substantially deteriorates for culturally and linguistically nuanced assessments, particularly for linguistic plausibility evaluation and in repetition detection in long dialogue sequences. Conversely, well-defined objective criteria, such as toxicity detection and problematic content identification, demonstrate robust inter-annotator alignment. Since our data is generated using benign prompts and topics, toxic or problematic content is expected to be very rare or nonexistent. These results are consistent with earlier research on the challenges of subjective versus objective evaluation and reveal shortcomings in current LLM-judges for assessing culturally sensitive content  (Watts et al., 2024). The distribution of scores across tasks and languages for both humans and LLM can be found in Appendix  A.2.
@@ 6 Downstream Tasks Evaluation
We selected two base models, Llama-3.1-8B and Phi4-14B, for fine-tuning experiments. These models were chosen based on their size (for feasibility of experiments given available resources) and reported multilingual capabilities (Grattafiori et al., 2024; Abdin et al., 2024). We used the Axolotl framework [^4] https://github.com/axolotl-ai-cloud/axolotl for all fine-tuning runs. Details regarding the hyperparameters and compute resources used can be found in Appendix A.4.
Baselines: We fine-tuned both Llama-3.1-8B and Phi4-14B on three high-quality, open-source IFT datasets that collectively cover nearly all the Indic languages included in our work: the Aya-Collection (Singh et al., 2024b), IndicAlign (Khan et al., 2024) and the Bactrian-X dataset (Li et al., 2023). To the best of our knowledge, these are the only open datasets that offer both broad language coverage and instruction-following format, and there is a lack of other large-scale open IFT datasets for Indian languages. The Bactrian-X dataset provides training samples for 10 out of our 13 target languages, excluding Assamese, Kannada and Punjabi. The Aya-Collection dataset includes all target languages except Punjabi. Since Aya-Collection contains millions of samples per language, we uniformly sub-sampled it to construct a balanced dataset of approximately 7 million samples which is comparable in scale to Updesh. For IndicAlign, the WordNet subset ($\sim$97M pairs) is disproportionately large, less diverse, and redundant, so we downsampled it to one instance per entry, yielding about 7.3M training pairs while retaining all other subsets.
Downstream Tasks: Our evaluation framework consists of three task categories to comprehensively assess model capabilities. Natural language understanding (NLU) tasks use multiple-choice questions to measure comprehension and reasoning through likelihood-based scoring. Natural language generation (NLG) tasks like translation and summarization, assess models’ ability to generate coherent and contextually appropriate outputs. For instruction-following evaluation, we translate established benchmarks (IFEval, IFBench) using GPT-4o (prompts in A.6), enabling assessment of both cross-lingual instruction-following (English responses to non-English instructions) and native language instruction-following. This design systematically identifies model strengths and weaknesses across diverse tasks, providing a holistic performance assessment. Dataset details are in Appendix A.5.
@@ 6.1 Results
Table 4: NLU task performance. All entries are accuracy (higher is better).
Model Setting MMLU MMLU-I MILU ARC-I BoolQ-I TVQA-I Bele INCL GMMLU
Llama Base 0.23 0.23 0.26 0.25 0.62 0.27 0.24 0.25 0.23
Bactrian 0.63 0.37 0.42 0.26 0.65 0.66 0.41 0.39 0.42
Indic-Align 0.60 0.34 0.42 0.27 0.71 0.59 0.32 0.34 0.39
Aya 0.59 0.33 0.39 0.28 0.76 0.51 0.47 0.34 0.37
Updesh 0.51 0.34 0.37 0.30 0.78 0.46 0.51 0.31 0.37
Phi4 Base 0.78 0.43 0.52 0.30 0.80 0.68 0.55 0.45 0.55
Bactrian 0.78 0.40 0.48 0.29 0.74 0.67 0.48 0.42 0.48
Indic-Align 0.77 0.45 0.51 0.32 0.76 0.62 0.52 0.45 0.54
Aya 0.78 0.45 0.53 0.33 0.73 0.72 0.61 0.48 0.53
Updesh 0.78 0.49 0.54 0.32 0.81 0.60 0.74 0.47 0.57
Performance on NLU Tasks: Table 4 shows the performance of both base models fine-tuned on all three IFT datasets, along with the performance of the base models on NLU tasks. Phi4-Updesh stands out as the most consistently strong configuration, attaining the best overall scores on MMLU-I, MILU, BoolQ-I, BeleBele, INCL, and GlobalMMLU. For the rest of the tasks, different models excel - in case of the Llama base model, Bactrian-X leads knowledge-intensive tasks, attaining the highest on MMLU and TVQA-I, while Updesh excels in instruction-style reasoning benchmarks (ARC-I, BoolQ-I) and performs best on reading comprehension tasks such as BeleBele. Phi4’s MMLU performance appears saturated across all fine-tuning baselines, suggesting a ceiling effect from pre-training knowledge or lack of coverage within all the training mixtures. Overall, these results demonstrate that different IFT datasets offer distinct advantages across various NLU capabilities.
Table 5: NLG task performance. All entries are ChrF scores (higher is better).
Model Setting XSum Flores En-XX Flores XX-En IN22-Conv-Doc
Llama Base 0.16 1.00 36.91 0.45
Bactrian 0.21 23.69 46.31 14.09
Indic-Align 9.26 25.18 2.29 10.98
Aya 0.24 22.60 0.53 6.73
Updesh 17.39 32.63 46.84 21.25
Phi4 Base 11.32 23.03 51.39 22.80
Bactrian 0.31 25.93 47.75 15.45
Indic-Align 0.27 27.48 1.77 13.82
Aya 0.37 26.51 7.63 14.01
Updesh 10.21 34.58 53.29 18.74
Performance on NLG Tasks: Table 5 demonstrates that Updesh is clearly superior to the other datasets on NLG tasks. Llama-Updesh emerges as the best configuration, achieving the highest performance across all generative tasks with substantial margins over all competing datasets as well as the base model. In case of Phi4, Updesh delivers the strongest translation performance, building effectively on the already capable base model. However, summarization results are more nuanced: the base model leads on XSum with Updesh being competitive, while the other fine-tuned models perform poorly. On long-context conversational translation (IN22-Conv-Doc) outperforms all fine-tuned variants, though it slightly lags behind the Phi4 Base model. Overall, we observe that the long-context generative training data in Updesh provides significant benefits for generative tasks.
Table 6: Instruction-following capabilities. All entries are accuracy (higher is better).
Model Setting mIFEval mIFBench
Native in English Native in English
Prompt Inst. Prompt Inst. Prompt Inst. Prompt Inst.
Llama Base 0.24 0.36 0.18 0.31 0.19 0.21 0.16 0.17
Bactrian 0.17 0.27 0.14 0.25 0.17 0.18 0.11 0.13
Indic-Align 0.16 0.24 0.14 0.23 0.17 0.19 0.16 0.17
Aya 0.13 0.22 0.11 0.21 0.18 0.19 0.18 0.19
Updesh 0.20 0.30 0.18 0.28 0.16 0.19 0.14 0.15
Phi4 Base 0.22 0.32 0.19 0.30 0.14 0.16 0.13 0.14
Bactrian 0.22 0.32 0.19 0.30 0.20 0.21 0.15 0.16
Indic-Align 0.16 0.25 0.15 0.24 0.15 0.17 0.14 0.16
Aya 0.15 0.25 0.14 0.24 0.20 0.21 0.19 0.20
Updesh 0.22 0.32 0.19 0.29 0.19 0.22 0.16 0.18
Performance on Instruction Following Tasks: Table 6 presents results on the GPT-4o translated multilingual versions of IFEval and IFBench, which evaluate fine-grained constraint adherence through highly specific requirements like character positioning, word counts, and formatting specifications - constraints notably absent in any training datasets. Native refers to prompts as well as expected output in regional language, while in English uses regional language prompts but expects English responses. Prompt-level evaluation measures whether all instructions in a prompt are followed, while instruction-level evaluation measures individual instruction adherence. Updesh emphasizes general instruction following involving task completion and output formatting, making these benchmarks an effective OOD evaluation for precise constraint adherence. For Llama, fine-tuning generally degrades performance, but Updesh shows smaller drops (mIFEval: 3-4 point lesser degradation than next best - Bactrian; mIFBench: comparable or better performance than other baselines, lags Aya on in English tasks). Phi4 exhibits different dynamics - Updesh achieves positive gains on mIFBench over the Base model (+4/+2 points Native; +1/+3 points English) while maintaining base-level performance on mIFEval, outperforming other fine-tuning approaches (Aya, Indic-Align) that show mixed or negative changes. Updesh provides robust instruction-following with minimal catastrophic forgetting compared to other baselines.
Language-wise trends: Figure 2 shows a clear trend: performance improves as language resources increase for both models and task types (NLU accuracy and NLG ChrF). We illustrate this with the resource taxonomy Joshi et al. (2020), ordered from 0 (lowest-resource) to 5 (high resource). The relative improvements with Updesh are most pronounced in low/mid-resource bands, effectively narrowing the gap to high-resource languages. The base model (without fine-tuning) is the weakest, with Bactrian-X and IndicAlign often competitive in the medium resource languages while Aya-Collection is more variable. Notably, generation benefits more from data quality: NLG improvements increase more dramatically as languages become lower-resourced, most clearly for Phi4-Updesh. Across language-resource tiers, Updesh is consistently strongest, especially at higher resource levels, reaching about $\sim$0.65 for Llama (NLU), $\sim$0.76 for Phi-4 (NLU) in terms of average accuracy, and the highest ChrF for Phi-4 (NLG).
Figure 2: NLU and NLG task performance grouped by language–resource class taxonomy from Joshi et al. (2020). Updesh yields the largest relative gains in low/mid-resource languages
6.2 Discussion
Quality evaluation: Our comprehensive quality evaluation ranges from basic filters to nuanced human evaluation. Experiments with LLM-based evaluators show that while LLM-judges do well on some metrics, they struggle with subtle multilingual and multicultural contexts. For example, they often misjudge how a language is typically spoken or written, and tend to lost track of persona consistency in longer dialogues. These issues motivate the need for scalable and reliable evaluation methods that go beyond surface-level checks.
Task-type sensitivity (NLU vs.NLG): We observe variation by task type - for instance, on NLU-style tasks (usually multiple-choice), no configuration is a clear winner across all datasets; performance depends on the base model and the specific benchmark. In contrast, on NLG-style tasks that score free-form generations, models trained on Updesh consistently perform strongly. This pattern suggests that the quality control and discourse control learned during training is transferred more directly to NLG than to choice-based NLU.
Evaluation protocol differences matter: Most NLU evaluations compute correctness via likelihood comparisons over (context, option) pairs, effectively testing discrimination among short alternatives. NLG evaluations, by design, require longer generations and measure faithfulness, adequacy, and fluency under open-ended prompting. The divergence in scoring objectives partially explains why a dataset that emphasizes long-form instruction-following (like Updesh) yields larger gains on NLG than on likelihood-based NLU. We find that the Bactrian dataset tends to perform best on NLU-style benchmarks, which can be attributed to format alignment and a larger proportion of MCQ or short-from prompts.
Distribution mismatches between train and test: Updesh contains longer-context samples and lacks short-text MCQ-style items; its task composition emphasizes multi-step, generative instruction-following. Even the MCQ subset which is part of the reasoning data in Updesh consists of long-form answers to MCQs, with close to no representation of short-text or direct answers. In contrast, many NLU benchmarks, however, are MCQ-heavy and short-context. This composition shift (task format, length, and response style) likely reduces the relative NLU gains for Updesh-trained models while amplifying NLG gains, showing that models generalize best to downstream tasks with evaluation formats that resemble their training distribution.
@@ 7 Conclusion
In this work, we examined synthetic data as a potential remedy for the scarcity of multilingual and multicultural resources. Through a comprehensive framework and systematic experiments across Indian languages, we identified effective strategies for data generation, quality assessment, and downstream evaluation, beyond English-centric norms. We built Updesh, a 9.5M 13-language IFT dataset using a culturally grounded, bottom-up pipeline. Our comprehensive evaluation spanning data generation, quality assessment (human and LLM-as-judge), and downstream tasks, revealed that synthetic data can potentially bridge resource gaps. However, no single recipe dominates across tasks and models, motivating approaches tailored to diverse linguistic and cultural contexts. We will release the Updesh dataset, evaluation protocols, and detailed analyses to enable transparent, reproducible progress.

### v2

@@ 4.2 Q-A for Generative Data
Inter-Annotator Agreement To assess the reliability of automated evaluation, we computed percentage agreement between the LLM-judge and human evaluators. Agreement varies notably across metrics (Figure 2), declining for culturally and linguistically nuanced aspects such as linguistic plausibility and repetition detection in long dialogue sequences. In contrast, objective criteria like toxicity detection and problematic content identification show strong alignment. As the data is derived from benign prompts and topics, toxic or problematic instances are expected to be rare or absent. These findings align with prior work on the difficulty of subjective versus objective evaluation and highlight persistent limitations of current LLM-judges in assessing culturally sensitive content (Watts et al., 2024). The complete distribution of scores across tasks and languages for both evaluations is provided in Appendix B.4.
Figure 3: Evaluation plots for models finetuned on Updesh vs existing datasets
Figure 4: Model Performance Landscape: NLU vs. NLG vs. Win Counts. The horizontal axis represents the average NLU (accuracy between 0-100), while the vertical axis represents the average NLG score (ChrF between 0-100). The size of each bubble corresponds to the number of specific datasets (12 tasks evaluated) where that the model outperformed all others. Updesh model (green) demonstrates the most dominant position with high generation scores and the largest number of task wins across both Llama and Phi settings.
Figure 5: NLG performance across 16 out-of-domain Indic languages on Flores. Updesh (red) achieves the highest average scores on both Llama-3-8B and Phi-4 architectures, outperforming standard baselines (Zero-shot) and comparable instruction-tuned models (Bactrian, Aya, IndicAlign).
@@ 5 Downstream Tasks Evaluation
We selected two base models, Llama-3.1-8B and Phi4-14B, for fine-tuning experiments. These models were chosen based on their size (for feasibility of experiments given available resources) and reported multilingual capabilities (Grattafiori et al., 2024; Abdin et al., 2024). We used the Axolotl framework [^3] https://github.com/axolotl-ai-cloud/axolotl for all fine-tuning runs. Details regarding the hyperparameters and compute resources used can be found in Appendix B.6.
Baselines
We fine-tuned both Llama-3.1-8B and Phi4-14B on three high-quality, open-source IFT datasets: Aya-Collection (Singh et al., 2024b), IndicAlign (Khan et al., 2024) and Bactrian-X (Li et al., 2023). To the best of our knowledge, these are the only open datasets that offer both broad language coverage and an instruction-following format. Bactrian-X covers 10 of our 13 languages (excluding Assamese, Kannada, Punjabi), while Aya-Collection includes all except Punjabi. Since Aya-Collection contains millions of samples per language, we uniformly sub-sampled it to 7M samples to create a balanced dataset comparable to Updesh. Similarly, for IndicAlign, since the WordNet subset ($\sim$97M pairs) is disproportionately large, less diverse, and redundant, we downsampled it to one instance per entry, yielding 7.3M training pairs when combined with its remaining subsets.
Downstream Tasks
Our evaluation framework consists of three task categories to comprehensively assess model capabilities. Natural language understanding (NLU) tasks use multiple-choice questions to measure comprehension and reasoning through likelihood-based scoring. Natural language generation (NLG) tasks, such as translation and summarization, assess models’ ability to generate coherent and contextually appropriate outputs. We augment standard dataset-NLU and NLG evaluations with comparative evaluations to understand model win rates. This design identifies model strengths and weaknesses across diverse tasks, providing a holistic performance assessment. Dataset details are in Table 2.
Dataset Source
NLU MMLU Indic (MMLU-I) SarvamAI (2025)
ARC Indic (ARC-I) SarvamAI (2025)
BoolQ Indic (BoolQ-I) SarvamAI (2025)
TriviaQA Indic (TVQA-I) SarvamAI (2025)
BeleBele (Bele) Bandarkar et al. (2024)
INCLUDE (INCL) Romanou et al. (2025)
Global MMLU (GMMLU) Singh et al. (2025)
NLG Extreme Summarization (Xsum) Singh et al. (2024a)
Flores English to Others (Flores EnXX) Goyal et al. (2022)
Flores Others to English (Flores XXEn)
IN22-Conv (IN22-Conv-Doc) - EnXX Gala et al. (2023) Gumma et al. (2025)
IN22-Conv (IN22-Conv-Doc) - XXEn
Table 2: Evaluation datasets
@@ 5.1 Results
Figure 3 presents a comparative performance analysis of the Llama and Phi-4 architectures across NLG and NLU tasks for a diverse set of 13 Indic languages. Broadly, we observe that models fine-tuned on the Updesh dataset (represented in green) consistently outperform existing baselines, including Bactrian-X, IndicAlign, and Aya-Collection. This performance advantage is particularly pronounced in NLG settings, where Updesh fine-tuned models demonstrate substantially higher performance across both high-resource languages like Hindi and Bengali, as well as lower-resource ones such as Assamese and Odia. Detailed results for the NLG tasks could be found in Table 12. In NLU tasks, Updesh maintains a competitive edge, often surpassing the strongest baselines highlighting the efficacy of the dataset in fostering robust multilingual understanding and generation capabilities.
Figure 4 illustrates the comparative performance of the models across three distinct dimensions: Understanding (NLU), Generation (NLG), and overall robustness (Win Counts). While the NLU scores (x-axis) show a competitive landscape with tight clustering among fine-tuned models, the Updesh setting (green bubble) distinguishes itself significantly in generation tasks, consistently achieving the highest placement in terms of scores. Crucially, the bubble size indicates that Updesh secures the highest number of ‘wins’ - 7 for Llama3 and 8 for Phi-4 -far surpassing other baselines like Bactrian and Aya. We observe that Updesh has more pronounced NLU performance gains in the Phi4 setting compared to the Llama settings.
Language-wise breakdowns and dataset-level averages are reported in Appendix B.8 (Tables 11 and 12).
Evaluation on Unseen Languages
To further understand robustness and cross-lingual transfer it is essential to test cross-lingual transfer to languages not seen in training dataset. In one of our NLG evaluation dataset(Flores) there are 16 languages which are not present in training data - Awadhi, Bhojpuri, Bodo, Chhattisgarhi, Garhwali, Haryanvi, Konkani, Maithili, Malvi, Manipuri, Marwari, Pashto, Rajasthani, Sanskrit, Santali and Tibetan.
As illustrated in Figure 5, Updesh consistently outperforms all the other baselines across different model architectures. On Llama-3-8B, it achieves a strong NLG score of 24.93, clearly surpassing all counterparts, and the trend holds for Phi-4, where Updesh (24.10) again leads across baselines. These results demonstrate that Updesh ’s curation strategy enables robust cross-lingual transfer, even to languages unseen during training. Exact results on all languages could be found in Table 15.
Setting NLU
Avg NLG
Avg Dataset
Wins
0-shot Baseline 28.64 12.69 0
Updesh - R + G 43.97 32.28 6
Updesh - G 48.48 19.73 7
$\Delta$ (vs R+G) 4.42 -12.54
Updesh - R 41.53 8.18 0
$\Delta$ (vs R+G) -2.43 -24.10
Llama
Setting NLU
Avg NLG
Avg Dataset
Wins
0-shot Baseline 56.21 28.58 3
Updesh - R + G 59.20 32.69 7
Updesh - G 59.18 32.04 3
$\Delta$ (vs R+G) -0.02 -0.64
Updesh - R 56.79 1.80 0
$\Delta$ (vs R+G) -2.41 -30.86
Phi-4
Table 3: Comparative performance of Updesh ablations. R and G denote the Reasoning and Generation subsets of the dataset, respectively. $\Delta$ values represent the performance difference compared to the full (R + G) model.
Ablations on dataset composition
To disentangle the relative contributions of the reasoning and generative subsets introduced in Sections 3, we conduct controlled ablations on the Updesh dataset. Specifically, we isolate the Reasoning (R) component - comprising translated subsets of the OrcaAgent-Instruct and OrcaMath datasets focused on multi-step reasoning and chain-of-thought supervision - and the Generative (G) component, which contains open-domain instruction-following and culturally grounded synthesis tasks derived from Indic Wikipedia content.
We train individual models on each subset (Updesh-R and Updesh-G) and compare them against the full combined dataset (Updesh-R+G) to quantify their respective effects on NLU and NLG performance. Results presented in Table 3 reveal a clear interaction between dataset composition and model behaviour. For both Llama and Phi-4, isolating the generative subset slightly improves NLU (e.g., +3.78+3.78 for Llama) but substantially reduces NLG quality, suggesting that reasoning data - being of translated nature is beneficial for translation performance. Conversely, training exclusively on reasoning data leads to a marked decline in both NLU and NLG metrics (37.06-37.06 for Phi-4 NLG). These results demonstrate that our bottom-up data generation approach is superior to naive translation.
Setting NLU
Avg NLG
Avg Dataset
Wins
0-shot Baseline 53.17 34.42 3
Updesh - R + G 56.61 38.73 3
Updesh - R + G (Seq len = 32K) 55.85 40.59 7
$\Delta$ (vs R+G) -0.75 1.86
Table 4: Ablations on training sequence length
Ablations on training sequence length
Additionally, given that many benchmark tasks contain shorter contexts, it is necessary to determine whether the long-context nature of the training data might have interference with short-context capabilities. Therefore, we evaluate a variant of the Updesh trained with a 32K444halved from the original 64K context window to examine the sensitivity of performance to sequence length constraints. Notably, we observe in 3, that the Phi-4 32K variant achieves the most balanced profile, securing the highest number of dataset wins ($7$) by maintaining strong NLU performance while further boosting NLG capabilities. This indicates that training sequence length is an important design consideration.
Rank Model ELO Wins Losses Ties Win Rate
1 Updesh-32K 1695.55 18760 6897 742 0.711
2 IndicAlign 1658.50 17447 7991 657 0.669
3 Updesh 1606.52 16168 8868 1111 0.618
4 Updesh Reasoning 1537.19 12237 12371 1792 0.464
5 Updesh Generative 1483.15 13752 11206 1442 0.521
6 Bactrian-X 1311.22 6862 17806 1554 0.262
7 aya 1207.86 2313 22400 1588 0.088
Table 5: Overall ELO Rankings comparing Updesh variants against baseline models for Phi-4. The Updesh-32K model demonstrates superior performance, outperforming both internal variants and external baselines like Bactrian-X and Aya-Collection
5.2 Comparative Cultural evaluations (ELO Rankings)
In Section 5, we evaluate the trained models and baselines on well-established academic benchmarks, which primarily measure performance on standardized tasks. However, such evaluations do not fully capture how useful these models are for real-world user queries spanning diverse domains, nor do they adequately reflect their helpfulness to everyday users in culturally grounded scenarios. To address this gap, it is essential to perform robust comparative evaluations in nuanced cultural contexts, focusing on non-academic, real-world questions posed by users. [^5] Due to the high evaluation cost, we perform this analysis only on Phi-4 checkpoints; we focus on Phi-4 since it consistently outperforms Llama in our other evaluations.
Thereby, we collect data following the collection process for Samiksha Hamna et al. (2025) and created a set of questions to assess the cultural relevance and helpfulness of LLM responses on practical, community-driven queries. Following the LLM-as-a-judge framework, we utilized GPT-4o to determine the superior response in pairwise comparisons between model checkpoints. Our evaluation encompasses 91,982 battles across seven models. To ensure statistical robustness, we evaluated all possible model pairings and randomized answer positions to mitigate positional bias, and calculate ELO ratings following the work in Boubdir et al. (2023). ELO ratings are calculated as the battles progress, where a higher rating indicates better comparative performance. This evaluation indicates that Updesh-32K significantly outperforms other baselines, with only IndicAlign showing competitive performance. Table 5 presents a comparative evaluation of model performance based on ELO ratings. The Updesh-32K model achieves the highest rating of 1696, establishing itself as the top-performing model in this evaluation set. It marginally outperforms Indic Align Cleaned (1659) and maintains a significant lead over the standard Updesh (1607). Notably, the Updesh models demonstrate substantial improvements over existing multilingual baselines, with the lead model scoring over 300 points higher than Bactrian-X Indic (1311) and nearly 500 points higher than the Aya-Collection Indic Sampled (1208), highlighting the efficacy of the Updesh dataset in providing more useful and grounded answers to India-centric, domain-specific queries.
Alignment between ELO Scores and Automated Benchmarks
As illustrated in Figure 11, there is a clear positive correlation between the ELO scores and the evaluated metrics across the board. Notably, the Updesh-32K (U32K) variant demonstrates superior performance, consistently clustering in the upper-right quadrant of all three plots. It achieves the highest ELO score ($\approx$ 1700) while simultaneously maintaining leading scores in NLU Average ($\approx$ 0.55), NLG Average ($\approx$ 35), and the total number of Dataset Wins. In contrast, baseline models such as Bactrian and Aya show mixed results; while Aya remains competitive in NLG tasks, it lags significantly behind U32K in the aggregate ELO ranking.
@@ 6 Conclusion
In this work, we examined synthetic data as a potential remedy for the scarcity of multilingual and multicultural resources. Through a comprehensive framework and systematic experiments across Indian languages, we identified effective strategies for data generation, quality assessment, and downstream evaluation, beyond English-centric norms. We built Updesh, a 9.5M 13-language IFT dataset using a culturally grounded, bottom-up pipeline. Our comprehensive evaluation spanning data generation, quality assessment (human, LLM-as-judge), and downstream tasks, revealed that synthetic data can potentially bridge resource gaps. Results show Updesh dominates across tasks and models, for both NLU and NLG settings. We will release the Updesh dataset, evaluation protocols, and detailed analyses to enable future research.
7 Limitations
Lack of Reliable Data Quality Estimation for multilingual synthetic data
Our comprehensive evaluation revealed that current LLM-based evaluators demonstrate variable reliability across quality dimensions, showing strong agreement with human judgments on objective metrics like toxicity detection (96-98%) but significantly lower concordance on nuanced aspects like fluency assessment and persona consistency (45-60%). This necessitates exercising caution when relying solely on LLM-based evaluations for quality estimation of multilingual synthetic data and highlights the need for more calibrated evaluators and robust evaluation frameworks.
Cultural Authenticity and Knowledge Base Limitations
Cultural authenticity remains challenging due to reliance on Wikipedia as the primary knowledge base, as many cultural customs and contextual nuances specific to under-represented Indian communities may lack sufficient documentation on Wikipedia, potentially resulting in incomplete cultural representations that might favor well-documented urban practices over rural or minority contexts.
Lack of Specialized Benchmarking
Limited benchmarks exist for evaluating cultural aspects and long-context/multi-turn capabilities in Indic languages, making systematic assessment of these crucial aspects difficult despite Updesh’s emphasis on these capabilities. While our framework covers general NLU, NLG specialized benchmarks for cultural knowledge and reasoning are needed to systematically evaluate and make progress. Although we conduct an comparative evaluation on real-world queries asked by human users, however we have used LLM evaluator for rating, while human evaluation would be a more fine-grained indicator of quality which we leave for future work.